### v1

# InstructPipe: Building Visual Programming Pipelines with Human Instructions

Abstract.
Visual programming provides beginner-level programmers with a coding-free experience to build their customized pipelines. Existing systems require users to build a pipeline entirely from scratch, implying that novice users need to set up and link appropriate nodes all by themselves, starting from a blank workspace. We present InstructPipe, an AI assistant that enables users to start prototyping machine learning (ML) pipelines with text instructions. We designed two LLM modules and a code interpreter to execute our solution. LLM modules generate pseudocode of a target pipeline, and the interpreter renders a pipeline in the node-graph editor for further human-AI collaboration. Technical evaluations reveal that InstructPipe reduces user interactions by 81.1% compared to traditional methods. Our user study (N=16) showed that InstructPipe empowers novice users to streamline their workflow in creating desired ML pipelines, reduce their learning curve, and spark innovative ideas with open-ended commands.
Visual Programming; Large Language Models; Visual Prototyping; Node-graph Editor; Graph Compiler; Low-code Development; Deep Neural Networks; Deep Learning; Visual Analytics
copyright: rightsretainedjournalyear: 2024copyright: rightsretainedconference: Under Review; booktitle: ;doi: XX.XXXX/XXXXXXXisbn: XXX-X-XXXX-XXXX-X/XX/XXccs: Computing methodologies Visual analyticsccs: Software and its engineering Visual languagesccs: Computing methodologies Machine learning
Figure 1. Existing visual programming requires users to select nodes, conceive pipeline structure, and then create the pipeline starting from an empty workspace. In contrast, InstructPipe facilitates users to build interactive pipelines directly from text instructions.
1. Introduction
A visual programming interface provides users with a node-graph editor to program. As opposed to writing code in a code editor, the node graph allows users to build a pipeline in a visual workspace with nodes and edges. This approach effectively reduces technical barriers for users to prototype creative applications. Advances in machine learning (ML) further stimulate growing interest in visual programming. Open-sourced ML libraries (e.g., TensorFlow (ten, 2015), PyTorch (Paszke et al., 2019), and Hugging Face (Wolf et al., 2019)) provide users with various encapsulated modules to accelerate AI project development and experimentation. Meanwhile, this also provides valuable protocols for visual programming developers to create systems for ML applications (Du et al., 2023a), where ML practitioners can interactively test “off-the-shelf” models on the node-graph editor. Recent foundational models like large language models (LLMs) (Vaswani et al., 2017; Brown et al., 2020; Anil et al., 2023) and findings on Chain-of-Thought (Wei et al., 2023) further stimulate a community-wise interest in visual programming (Wu et al., 2022a, b; Du et al., 2023b), which provides users interactive experiences to explore AI chains.
Despite the development of visual programming platforms in various domains, we observed that existing systems share one similar characteristic: users usually start a creative process in the workspace “from scratch”. This implies that users need to 1) select nodes, 2) ideate the pipeline structure, and finally, 3) connect nodes from a completely empty workspace. For users who are unfamiliar with a particular visual programming platform, such processes can easily overwhelm them, degrading their overall programming experience.
Similar issues also exist when users write programs using text-based editors (there exist many built-in functions in a particular programming language and multiple variables in a program), but advances in LLM assistants show that such challenges can be effectively reduced. For example, GitHub Copilot (GitHub, 2023) makes it possible for users to generate code by simply describing users’ requirements in natural language. Even though the generated code is not absolutely correct, the AI assistance usually finishes a large portion of the task, and programmers may only need to make a few edits to achieve a correct result. To this end, we raise the following questions that motivate our work: How can we build such an assistant to benefit visual programming users?
In this paper, we built InstructPipe, an AI assistant for visual programming users to generate a pipeline through natural language instructions (Figure 1). We implemented InstructPipe on Visual Blocks (Du et al., 2023a), a visual programming system for prototyping ML pipelines. One major technical challenge in implementing InstructPipe lies in the lack of visual programming data. Different from the standard approach to building a copilot for the text-based editor (i.e., training LLMs using large-scale text-based programs online), it is impractical to collect sufficient data for a particular visual programming platform. We addressed this issue by decomposing the generation process into three steps. Using two separate LLM modules, the system first scopes the potentially useful nodes and then generates pseudocode for a target pipeline. InstructPipe then compiles the pseudocode and renders the pipeline on the node-graph editor to facilitate further user interaction. Our technical evaluation suggests that InstructPipe reduces user interactions by $81.1\%$ compared to building pipelines from scratch. Our system evaluation with 16 participants demonstrated that InstructPipe significantly reduced users’ workload in their creative process. Qualitative results further reveal that InstructPipe effectively supports novices’ on-boarding experience of visual programming systems and allows them to easily prototype a concept for various purposes. As one pioneering work on visual programming copilot, we also observed new challenges caused by humans’ cognitive characteristics and proposed future technical directions toward a next-level, open-ended AI prototyping assistant.
In summary, this work offers the following contributions:
  1. (1)
    InstructPipe, an AI assistant that enables users to build ML pipelines from human instructions,
  2. (2)
    System design and technical development of InstructPipe, which includes two LLM modules and a code interpreter that generate codes for a visual programming pipeline, compile the code, and render the pipeline in an interactive node-graph editor,
  3. (3)
    A technical and a user evaluations that demonstrate effectiveness of InstructPipe, and the corresponding findings that reveal new challenges for the HCI community.
2. Related Work
2.1. Visual Programming
A computer program defines the operation of computer systems. However, “the program given to a computer for solving a problem need not be in a written format” (Sutherland, 1966). This future-looking statement, dating back to the 1960s, inspired several generations of researchers to design and build visual programming systems.
Today, visual programming systems (e.g., LabView (Kodosky, 2020), Unity Graph Editor (Unity, 2023), PromptChainer (Wu et al., 2022a), and Visual Blocks (Du et al., 2023a)) typically feature a node graph editor, providing users with a visual workspace to “write” their program by “building blocks” (Hewett, 2005; Sramek et al., 2023). Recent work further explored the application of visual programming in education (Jiang, 2023; Kovalkov et al., 2020), authoring support (Zhang et al., 2023b; Zhang and Oney, 2020), and robotics (Datta et al., 2012; Huang et al., 2016; Huang and Cakmak, 2017). For example, Zhang et al. (Zhang et al., 2023b) connected the visual programming tool to the concept of teaching by demonstration (Myers, 1986; Li and Landay, 2005; Zhou and Yatani, 2022), allowing users to rapidly customize AR effects in video creation. FlowMatic (Zhang and Oney, 2020) extended traditional visual programming interfaces into 3D virtual environments, providing users with immersive authoring experiences.
More recently, findings on LLM Chains (Wu et al., 2022b) and Chain-of-Thought (Wei et al., 2023) further stimulated researchers to build visual programming tools that chain LLM modules. Developers want to explore various ways to chain an LLM module for various applications. In such a scenario, visual programming provides a great platform for users to focus on high-level exploration. Example research work and industrial products include PromptChainer (Wu et al., 2022a), LangFlow (LangFlow, 2023), and ComfyUI (ComfyUI, 2023).
This work offers technical contributions of implementing and evaluating an AI assistant that supports users to prototype an AI visual programming pipeline from text-based instructions. Compared to a mainstream user workflow in which people begin their creation from scratch, InstructPipe provides a new experience for beginner-level users to easily prototype a pipeline.
2.2. LLMs and Their Applications in Interactive Systems
Early multimodal ML work uses language models to solve Visual Question Answering (VQA) (Andreas et al., 2017; Antol et al., 2015), yet these solutions are limited to simple questions and cannot perform effective reasoning and problem solving (Ma et al., 2023). LLMs revolutionized AI’s reasoning capability (Wei et al., 2023; Zhou et al., 2022) in language, which motivate researchers to build LLM applications in various domains beyond NLP (Surís et al., 2023; Singh et al., 2023; Park et al., 2023). For instance, LLMs augmented the perceptual and planning intelligence in robotics (Singh et al., 2023), supported autonomous driving (Liang et al., 2022), and assisted clinical processes in medical science (Kung et al., 2023; Singhal et al., 2022).
The advances in LLMs empower recent interactive systems (Peng et al., 2023; Park et al., 2022) with enhanced machine intelligence. Recent research leverages LLMs to edit visualization (Wang et al., 2023a; Shen et al., 2023), receive AI explanation (Wu et al., 2021), facilitate communications (Liu et al., 2023), understand user interface (Wang et al., 2023b), and study simulated social behaviors (Park et al., 2023). The revolution also motivated HCI researchers to design new interfaces for LLMs (Jiang et al., 2023; Suh et al., 2023). Graphologue (Jiang et al., 2023) augmented LLM responses by displaying interactive diagrams along with the response text, which depicts the semantic logic in a paragraph. Sensecape (Suh et al., 2023) provides users with a workspace to explore long LLM responses in a hierarchical structure.
InstructPipe focuses on utilizing LLMs for generating pipelines in visual programming with human instructions (Hu et al., 2023; Surís et al., 2023). This work shares a similar vision with Prompt2Model (Viswanathan et al., 2023) and VisProg (Gupta and Kembhavi, 2023). Prompt2Model (Viswanathan et al., 2023) finetunes a BERT model (Devlin et al., 2019) using data generated by instructions. VisProg (Gupta and Kembhavi, 2023) produces Python code from instructions with task-dependent few-shot prompting. However, prior arts focus on full AI automation and lack an interactive workspace for human-AI collaboration. In contrast, InstructPipe generates and compiles a pipeline (without finetuning), while rendering the pipeline in a visual programming interface, facilitating an interactive, collaborative, and explainable workflow.
3. InstructPipe
((a)) InstructPipe’s instruction dialog.
((b)) InstructPipe’s visual programming interface.
Figure 2. The user interface of InstructPipe. The user can first click on the “InstructPipe” button on the top-right corner of the interface in (b). A dialog will appear, and the user can input the instruction and select a category tag. InstructPipe then renders a pipeline on (b), in which the user can interactively explore and revise.
InstructPipe is an AI assistant that enables users to generate a visual programming pipeline by simply providing text-based instructions. We implemented InstructPipe on Visual Blocks (Du et al., 2023a), a visual programming system for prototyping ML pipelines.
3.1. User Workflow
To generate a pipeline, users can first click the “InstructPipe” button on the top-right corner of the interface (1(b)). The system then activates a simple dialog (1(a)) in which users can 1) provide a description of their target pipeline and 2) tag the pipeline. The tag can be “language”, “visual”, or “multimodal”. After users click the “Submit” button in the dialog, InstructPipe renders a visual programming pipeline on the node-graph editor. Based on the result, users can further refine the pipeline in the visual programming interface.
3.2. Primitive Nodes
Figure 3. An overview of major primitive processor nodes supported by InstructPipe. 17 of the 20 processor nodes are categorized by the type of accepted I/O into a $3\times 3$ matrix. Note that “PaLM” represents two nodes in InstructPipe, i.e., a text generation model and a chat model of PaLM (Anil et al., 2023).
InstructPipe supports 27 primitive nodes in Visual Blocks, including seven I/O nodes (e.g., live camera and markdown viewer) and 20 processor nodes. Appendix A provides a description for each of the 27 nodes. Compared to related work (Gupta and Kembhavi, 2023; Surís et al., 2023) that automates ad hoc ML inferences in specific use scenarios, we aim at an open-ended use case with diverse primitive nodes. Figure 3 visualizes 17 (out of 20 in total) processor nodes according to the I/Os of the nodes. The remaining three nodes are “Google Sheet” (which takes a Google Sheet URL as input and outputs the sheet data), “image_mixer” (which combines multiple images), and “virtual_sticker” (which casts a sticker on a person’s face on a live camera). “Metadata” in Figure 3 indicates intermediate data used in ML pipelines, which can be a segmentation mask that describes an input image, or a URL that describes a target news that users want to read. As shown in the matrix, InstructPipe contains a wide range of nodes that supports the creation of complex ML pipelines.
4. Pipeline Generation from Instructions
Figure 4. Workflow of InstructPipe. First, users describe their desired pipeline in natural language and designate it with a language, image, or multi-modal tag. InstructPipe then feeds user instructions into a node selector to identify a relevant set of nodes. Subsequently, both the instructions and the relevant nodes with their description are input into a code writer to produce pseudocode. Finally, a code interpreter parses the pseudocode, rectifies errors, and compiles a JSON-formatted pipeline, allowing users to refine and interact with it further within Visual Blocks’s node-graph editor.
InstructPipe leverages LLMs to generate visual programming pipelines from text instructions. There exist two prevailing approaches for customizing LLMs: 1) fine-tuning LLMs (Brooks et al., 2023; Liu et al., 2023), and 2) few-shot prompting (Park et al., 2023; Dong et al., 2023). For our task, fine-tuning LLMs requires a substantial volume of annotated data comprising pairs of pipelines and prompts. Additionally, a growing list of nodes poses new challenges to scaling this approach with new data annotations. In comparison, few-shot prompting can seem more practical (Gupta and Kembhavi, 2023; Wei et al., 2023; Yao et al., 2023), but it is challenging to design an efficient prompt that fits within a reasonable number of tokens. The configuration file of the 27 nodes alone includes 8.2k tokens. Moreover, because of the combinatorial explosion of the 27 nodes in the system, it is not clear how many prompt examples are needed and how we can construct in-context pipeline examples.
To this end, we implement InstructPipe with a two-stage LLM refinement prompting strategy, followed by a pseudocode interpretation step to render a pipeline.
((a)) Pipeline.
((b)) pseudocode.
Figure 5. A pair example of pipeline and pseudocode. In the first line of code under “processor”, pali_1_out, pali_1, pali and image=input_image_1, prompt=input_text_1 represents output variable id, node id, node type, and node arguments, respectively.
Figure 4 illustrates the high-level workflow of the InstructPipe implementation. InstructPipe includes two LLM modules (highlighted in red): 1) a Node Selector (section 4.2) and 2) a Code Writer (section 4.3). Given a user instruction and a pipeline tag, we first devise the Node Selector to identify a list of potential nodes that would be used according to the instruction. In the Node Selector, we prompt the LLM with a very brief description of each node, aiming to filter out unrelated nodes for a target pipeline. The selected nodes and the original user input (the prompt and the tag) are then fed into the Code Writer, which generates pseudocode for the desired pipeline. In Code Writer, we provide the LLM with detailed description and examples of each selected nodes to ensure LLMs have a thorough understanding of each candidate node. Finally, we employ a Code Interpreter to parse the pseudocode and render a visual programming pipeline for the user to interact with.
4.1. Pipeline Representation
The Visual Blocks system represents a pipeline as a Directed Acyclic Graph (DAG) in JSON format [^1] Pipeline in JSON files: see supplementary materials for examples.. To compress the verbose JSON file, InstructPipe represents pipelines as pseudocode (Gupta and Kembhavi, 2023; Surís et al., 2023), which can be further compiled into a JSON-formatted pipeline. The pseudocode representation is highly token-efficient. Figure 5 shows an example in which the pseudo compresses the JSON-based pipeline representation (2.8k tokens) into a 123-token pseudocode representation. The efficiency does come with a cost: it loses some fine-grained annotations of each node like property values (e.g., layout of the nodes, parameters of a segmentation model, and the degree of blurring parameters in an “image processor” node). InstructPipe leaves the task of finetuning these parameters to the user and focuses on generating the graphic structure of a visual programming pipeline.
Figure 5 provides an example of a pipeline (4(a)) and its corresponding pseudocode (4(b)). The syntax design is inspired by TypeScript. The structure is inspired by how academic papers present pseudocode (Zhang et al., 2021): an algorithm block typically starts with specifying the input/output and then explains the intermediate processor. We highlight the first line under the processor module (i.e., the operation of the PaLI node) in 4(b) in four different colors, representing four different components in the programming language. “pali_1_out” represents the output variable name of the node. “pali_1” is the unique ID of the node. The green symbol after the colon, i.e., “pali”, specifies the type of the node. In this example, the node with the ID of “$pali\_1$” is a “pali” node. The rest part in the bracket, i.e., “image=input_image_1, prompt=input_text_1”, defines the arguments of the nodes. In the input pseudocode, we do not annotate the output variable (i.e., there are no red colors in the highlighted line under the input module) because all the input nodes only export one value, and the output variable name is automatically annotated as the same symbol as the node id (i.e., “input_text_1”). Note that InstructPipe generates the text input (i.e., the property value) in the “input text” node. Therefore, the argument in “text=“caption this image in detail”” does not indicate that the “input_text” node accepts input edges, but accepts the node property input.
[Uncaptioned image] Figure 6. The prompt structure for the Node Selection module. Each node description is formated as ”{node types}: {short descriptions of the nodes}; {recommended node(s)}”. The node recommendation is optional.
[Uncaptioned image] Figure 7. The prompt structure for the Code Writer module. Detailed node configurations, see supplementary materials for examples, are listed in the highlighted region.
4.2. Node Selector
Node Selector aims to filter unrelated nodes by providing an LLM with a short description of each node. Figure 6 shows the prompt we use in Node Selector. Followed by a general task description and guidelines, we list all the node types with a short description that explains the function of the node. Several nodes come with recommendation(s) when the users interact with Visual Blocks, and we also include such node recommendations information in the prompt. The main intuition of this prompt design is based on how the existing open-source libraries (e.g., Numpy (Harris et al., 2020)) present a high-level overview of all functions [^2] See an example in the following link: https://numpy.org/doc/1.25/reference/routines.array-manipulation.html. The documentation typically presents a list of supported functions (in each category), followed by a short description so that developers can quickly find their desired functions. At the end of the prompt, we provide a list of Q&A as few-shot examples to support the LLM to learn and adapt to the context of the task.
4.3. Code Writer
With a pool of selected nodes, the Code Writer module can write pseudocode for rendering a target pipeline. Figure 7 shows the structure of the prompt utilized in this LLM module. Similar to section 4.2, the prompt starts with a general introduction and several guidelines. The major difference in the prompt design in this stage lies in the granularity of each node introduction. We provide a detailed configuration for each selected node with additional information, including 1) input data types, 2) output data types, and 3) an example, represented in pseudocodes, showing how this node connects to other nodes. We put a detailed explanation of the full node configuration in the supplementary materials. Similar to the previous LLM module (section 4.2), the prompt design here is also inspired by the documentation of several popular code libraries. Specifically, we gain inspiration from low-level function-specific documentation [^3] See an example in numpy.shape: https://numpy.org/doc/1.25/reference/generated/numpy.shape.html#numpy-shape, which typically includes a detailed description and data types in the input/output, followed by one or more examples of how developers can use this function with a few lines of codes.
The prompt also includes a list of Q&A as few-shot examples. However, providing few-shot examples in this stage is non-trivial. The reason lies in the dynamics of the node selection pool: a combination of all the nodes causes many possible selections, and it is impossible to design a list of few-shot examples for each possibility. Therefore, we only created an example list for each pipeline tag (i.e., “language”, “visual”, and “multimodal”) and intended to utilize these few-shot examples to cover most of the use cases. The intuition behind this design is: the in-context examples serve to list possible creative ideas of the pipelines in each tag, which can be condensed into few-shot examples rather than traversing all the possible combinations. This implies that the in-context examples may include nodes that are not selected in the prompts. According to our preliminary tests, such out-of-scope nodes cause negative effects on the generation results: LLMs tend to also “invent” new nodes that do not exist in our system, causing failure in node rendering. Note that we also observed this issue when combining the prompts used in (Gupta and Kembhavi, 2023). We eliminate this issue by using the prompt contents between the node configurations and the in-context examples (i.e., the contents start with “the following is a full list of …” in Figure 7). This aims to inform LLMs of the exceptions for invention.
((a)) Before layout optimization.
((b)) After layout optimization.
Figure 8. A comparison of the same generated pipeline before and after layout optimization. (a): Each node is assigned with a default property value when convert pseudocode into a JSON file for rendering the pipeline. Such default values will cause sub-optimal visual effects. (b): Our layout optimization process re-arranges the layout for better presentation of the pipeline.
4.4. Code Interpreter
Finally, InstructPipe employs a code interpreter to parse the generated pseudocode and compile a JSON-formatted pipeline with an automatic layout. We delineate the graph compilation and rendering procedure below:
  1. (1)
    Lexical Analysis: InstructPipe first tokenizes each line of the pseudocode is into output variable id, node id, node type, and node arguments (section 4.1).
  2. (2)
    Graph Generation: The tokenized results allow us to build a graph structure in a node-graph editor that connects each node as specified by the pseudocode. We then generate JSON-formatted code, which Visual Blocks uses to render the pipeline in the node-graph editor. Note that the JSON code includes far more parameters than those used in InstructPipe for defining the graphic structure. InstructPipe fills texts in the “Input Text” node based on LLM-generated pseudocode and uses default values for the rest nodes. For example, by default, the temperature and the max output tokens for the PaLM node are set to 0.5 and 256, respectively. If users are not satisfied with the default values, they can interactively adjust the parameters in the node-graph editor.
  3. (3)
    Graph Rendering and Optimization: A problem with employing the default parameters is that it results in a chaotic distribution of nodes in the node-graph editor: each node is located in a predefined position, and the edges go across the workspace without a reasonable arrangement. Therefore, InstructPipe traverses the graph with a breadth-first search (BFS) and arranges the nodes based on their depth values in the DAG. As shown in Figure 8, such post-processing can effectively enhance the visualization of a generated pipeline.
Our exploration shows that LLMs may not always generate accurate code as expected. One typical issue is that Code Writer tends to invent illegal nodes out of our library of 27 nodes. To address this, InstructPipe disposes such lines that leverage illegal nodes. If legal nodes use the output of such an illegal node, InstructPipe discards the edge connection between them to prevent bugs when running the pipeline. Another common issue is that the generated code may not have a correct order. For example, the generated code may be “Line 1; Line 3; Line 4; Line 2;”, in which “Line 1 – 4” represents four lines of code, and the correct order should be from “Line 1” to “Line 4”. Such ordering issues would cause an input value of the current line to become undefined, because its definition is mistakenly placed afterward. We address this issue by stacking the line that is not ready for execution, similar to graph construction by traversing an adjacency list. The Code Interpreter runs in a loop to interpret the code until all legal lines in the stack become ready.
5. Technical Evaluation
Evaluating our system is challenging. Ideally, the evaluation set should cover 1) diverse visual programming pipelines (i.e., the pipeline factor) and 2) a variety of instructions created by different individuals for each pipeline (i.e., the human factor). However, such a combination requires a significant number of user evaluation sessions, which is impractical to deploy. For example, if an evaluation set includes 50 pipelines and 20 instructions for each pipeline from 20 different participants, it requires 1,000 user study sessions. Asking participants to attend multiple pipeline creation sessions seems efficient, but it introduces new learning transfer effects to the study. The learn effects usually require additional user study design with counterbalancing, which eventually increases the total number of required sessions.
To address this issue, we decouple the two factors that cause variations by decomposing the whole evaluation into two evaluation sessions. The major benefit of this design is that it allows us to deploy the evaluation tasks with a reasonable workload while maintaining the rigorousness of the whole evaluation. In our first evaluation session (i.e., technical evaluation), we focus on understanding the technical performance of InstructPipe among a variety of pipelines. In the second evaluation session (i.e., user evaluation), we control the pipeline factor and examine the human factor by recruiting different participants for prototyping controlled pipelines. The first technical evaluation provides crucial data on the “accuracy” of InstructPipe, which guides us to select such representative pipelines with as little bias as possible. More importantly, the user evaluation further allows us to understand how user experiences are affected under two conditions: with and without InstructPipe.
In this section, we focus on explaining the first step of our evaluation, the technical evaluation, and the following user evaluation is covered in the next section.
5.1. Data Collection
To conduct the technical evaluation, we first organized a two-day hybrid workshop with 23 participants, aiming to collect pipelines that Visual Blocks users build for their creative usage. Six attendees claimed that they had prior experience in using Visual Blocks. At the beginning of the workshop, we gave a 15-minute tutorial walking the participants through the nodes and the pipeline-building process using Visual Blocks. After the tutorial, attendees created pipelines independently. We required the participants to export the JSON file from Visual Blocks and upload it to our Google Drive every time they finished a pipeline. We also emphasized that they should caption their pipelines when they export the files. Such data pairs (pipeline/caption) by the participants constitute important raw data we use for the technical evaluation.
Note that Visual Blocks includes more primitive nodes than the 27 nodes covered by InstructPipe. The workshop is an open-ended creation process, in which participants are free to any node available in Visual Blocks. The InstructPipe feature was not available in the workshop.
5.2. Data Post-Processing
((a)) Search news from Google, summarize it and then conduct fact check. Input: a keyword for Google Search; Output: a summarization of the news and a fact-check result.
((b)) Generating an emoji from a photo. Input: a photo uploaded by the user; Output: an emoji generated from the photo.
((c)) Turning a tiger into a cat. Input: an image of a tiger; Output: an image of a cat.
Figure 9. Example pipelines participants built in the workshops.
After the workshops, one author carefully examined each collected pipeline and found several critical issues in the raw data:
  • Incomplete pipelines. There exist pipelines uploaded by the participants that were incomplete.
  • Isolated graphs. There exist pipelines that include at least one isolated subgraph. The isolated subgraph, as opposed to the main graph, is defined as a graph (or a node) that has no connection to the main graph in the pipeline that provides the main functionality of the pipeline. The “Image viewer” node on the bottom-left corner of 7(b) We observed that some participants typically would like to explore the system by working on a separate sub-space. While we acknowledge its usefulness, leaving such “redundant” graphs in the raw data for the evaluation would cause issues when we calculate the number of user interactions (i.e., the metric used in the evaluation that will be defined in the next subsection).
  • Low-quality captions. While we explicitly required the participants to write descriptive captions, we found some captions written by the participants were either empty or low-quality (e.g., “newsletter”, “image editing” and “[participant name]-demo”).
The observation motivated us to post-process the raw data to present more rigorous evaluation results. We first removed incomplete pipelines and the isolated graphs in each pipeline (if there are any). Although the low-quality captions may reflect real-world scenarios on how the users may use our system, such a hypothesis cannot be officially validated. In our data collection process, participants are situated in the scenarios of writing captions instead of instructing an AI assistant. When users generate instruct pipelines, they may be aware of the importance of writing good instructions, and thus, such low-quality captions would not be as usual. More importantly, the technical evaluation focuses on considering the variation of the pipeline, instead of text instructions. In the next user evaluation (section 6), we will explicitly situate a new group of participants in instructing the assistant and observe the user behaviors and the variation of human instructions on the same pipelines. Therefore, we were motivated to enhance the quality of the captions for fair technical evaluations. Two authors individually annotated the caption of each pipeline separately by referring to the original captions and pipelines authored by the participants. It is important to note that we finished the workshop and the data annotation task before we completed the system implementation. The two authors had no experience using InstructPipe before completing the annotation. We believed this process could effectively enhance the quality of the captions while maintain fairness of the technical evaluation.
As we clarified in section 5.1, the workshop is designed to be an open-ended creation process. This indicates that the dataset inevitably includes out-of-scope nodes like “custom scripts” (in which the participants write code to process the input data and return custom outputs; see 8(b) for an example) and “TFLite model runner” (which call a custom TensorFlow model with a URL input of the model in the TF-Hub). As we explained in section 4.1, InstructPipe focuses on generating the graphic structure of a pipeline without considering the property value within each node. Different from the majority of nodes in Visual Blocks, nodes like “custom scripts” and “TFLite model runner” derive their intrinsic functionality from the property values. Without this information, the entire pipeline’s rationale is obscured. Therefore, we removed pipelines that contained “out-of-scope” nodes in our technical evaluation.
The final 48 pipelines (out of 64 pipelines) are comprised of 23 language pipelines, seven visual pipelines, and 18 multi-modal pipelines. Figure 9 shows three pipelines created by the participants. 8(b) is an example of the pipelines that include out-of-scope nodes, and therefore is not included in the final 48 pipelines. In the technical evaluation, we ran our generation algorithm on the pipeline captions six times (three times for each caption $\times$ two captions from two authors for each pipelines) and evaluated the generation results using the metric that will be introduced below.
5.3. Metric: The Number of User Interactions
Evaluating a generated pipeline is more complex than other tasks that have a universally recognized definition of accuracy. In this project, intuitively, an accurate generation implies that users only need to do very few edits (or even no edits) based on the generated result. This intuition inspires our definition of “the number of user interactions” as the core metric used in our evaluations:
The Number of User Interactions is defined as the minimal number of user interactions needed to complete the pipeline from a generated pipeline.
Note that there are countless ways to modify a generated pipeline toward a complete pipeline in practice. Nevertheless, the minimal number of user interactions, representing the “smartest” way(s) to make a modification, is deterministic, and this is an objective metric that can fairly reveal how many minimal efforts users need to spend to achieve their goal. A pipeline is considered complete when it satisfies the given instruction. We calculate the number of interactions from the sum of two events: 1) adding or deleting a node and 2) adding or deleting an edge between nodes. Note that when a node with edges connected is deleted, our system will auto-remove these edges. In such instances, we only register one interaction for the node deletion.
Our definition of the number of user interactions has two important implications. First, a complete pipeline after user interaction does not need to be the same as the corresponding pipeline in the dataset. As long as it fulfills the task described in the caption, we consider the pipeline complete. Second, our definition does not consider interactions of modifying property values of a node, e.g., typing in a text box or selecting a value in a drop-down box. We argue that such interactions are highly node-dependent and are hard to quantify objectively. More importantly, as we explain in section 4.1, the generation of property values is out of the scope of this work.
In the technical evaluation with various pipelines, it is unfair to report an averaged absolute value of user interactions because the complexity of the pipelines varies dramatically. For instance, the user may need three edits based on a generated result to complete a large pipeline that requires 20 edits from scratch. In another pipeline, the user also needs to do three edits starting from the generated result, but the whole pipeline only takes three edits to finish. Averaging these absolute values does not provide reasonable insights into how accurate the generation is. Therefore, we reported an averaged ratio of user interactions required to complete a pipeline “from our generated pipeline” to that “from scratch” as our target metric in the technical evaluation.
Table 1. The ratio of human interactions in the technical evaluation. Results are reported as mean $\pm$ standard deviation.
Overall Language Visual Multimodal
$\mathbf{18.9\pm 20.3}\textbf{\%}$ $17.4\pm 20.6\%$ $17.6\pm 23.7\%$ $20.8\pm 16.0\%$
5.4. Results
Table 1 summarizes the results of the technical evaluation. Compared to building a pipeline from scratch, InstructPipe allows the user to finish a pipeline with 18.9% of the user interactions (as defined in section 5.3), demonstrating the effectiveness of the InstructPipe support. Seven generated pipelines directly satisfied with instructions without user interactions in all six trials, and 38 generated pipelines completed at least once in any of the six trials.
6. User Evaluation
While the technical evaluation demonstrates the accuracy of InstructPipe among various real pipelines created by participants, it is still unclear what is the actual user experience when real users go through the entire system workflow. Additionally, there is a lack of subject variation on a controlled group of pipelines in the technical evaluation. Therefore, we conducted an in-person user study of InstructPipe with another group of participants, aiming to provide more insights into our system performance. The study recruitment was in accordance with the ethics board of our institution. We obtained participant consent before the study began.
6.1. Study Design
Figure 10. A flow diagram of the user study. After a training session, participants completed the two tasks in each condition in the sequence determined by the counterbalancing protocol.
In the user evaluation, we aimed to investigate how the interface condition (with InstructPipe and without InstructPipe; the independent variable) affects the user experience and behaviors (the dependent variable). We will refer to these two interface conditions as “InstructPipe” and “Visual Blocks” in the following content. Figure 10 visualizes the complete study flow. In each condition, participants completed the two pipelines with counterbalance (referred to as Task 1 and Task 2 in Figure 10).
We carefully designed this study to ensure a fair study that can be completed with reasonable efforts. The following elaborates how we made two important decisions that affected the rigorousness of our study:
  • Two controlled pipelines. Our user evaluation focuses on two controlled pipelines. While we acknowledge more pipelines (e.g., four, six, or even more, instead of two in our design) can enhance the solidness of the study, such design also dramatically increases the required groups of the user study session. For example, counterbalancing four controlled pipelines with two interface conditions results in $4!\times 2=48$ study orders. If there are four participants for each unique order, the whole study requires $48\times 4=192$ participants, which is far more than a standard number in HCI system papers (i.e., recruiting 10 - 20 participants for evaluation). We believe two pipelines fit the best in our case since they require $2!\times 2\times 4=16$ participants.
  • Pipeline selection. With the limited number of pipelines we can choose in the study, it becomes highly important how we select the two pipelines. There are two critical factors we considered in the pipeline selection process: representativeness and diversity. The representativeness implies that the selected pipelines should represent the averaged performance of InstructPipe. The diversity further suggests that the selected pipelines should provide various experiences to simulate the actual use scenarios in which InstructPipe may work well in some pipelines but not always. By following this guideline, we first selected four candidates, and the final decision was made after a pilot study with one participant to test the level of pipeline difficulty. The two resultant pipelines are composed of eight nodes with seven edges and six nodes with six edges, respectively. Using the instructions from two authors, the averaged ratio of human interactions in these two pipelines are $27.8\%$ and $5.2\%$, respectively. See Appendix C for more details of the pipelines.
6.2. Procedure
The study starts with a 10-15 minute hands-on training of both conditions. The training included 1) all the Visual Blocks interactions needed to complete the subsequent steps of the experiment and 2) all the nodes that participants need to use for creating the pipelines they are assigned in the main sessions. Participants were also encouraged to try building a pipeline independently and to ask questions.
After the training, participants progressed to an unmoderated session where they were asked to build pipelines under the given conditions. We verbally described the pipelines to participants as below, and participants are not allowed to read our scripts:
  • Text-based pipeline: get the latest news about New York using Google Search and compile a high-level summary of one of the results.
  • Real-time multimodal pipeline: create a virtual sunglasses try-on experience using your web camera.
During the task, participants were allowed to consult with us for technical help. If participants were unable to make progress, we provided hints. We provided such support to ensure that each participant spent a reasonable amount of time on each task and had sufficient time for the following tasks. As shown in the results (section 6.5.2), we provided far more support in the “Visual Blocks” condition. Therefore, we argued that this is a fair moderation in the study because it favors our comparative conditions (i.e., it makes it easier for users to finish the pipeline under the “Visual Blocks” condition). As an optional extension to the study, eight participants were offered an open-ended pipeline creation, where participants prototype their own ideas with InstructPipe. This optional section was offered based on the progress of the participant in the previous sections and time constraints so that the study duration was controlled within the time we guaranteed in our recruitment process.
After trying all pipeline-condition combinations, participants answered open-ended questions in a semi-structured interview. The interview script is available in the appendix §D. Participants provided their general outlook of each condition, listed pros and cons, identified potential future use cases, and critiqued the user interface for future improvements.
In total, participants spent 55-65 minutes in the study.
6.3. Participants
We recruited 16 participants from an internal participant pool at Google. A diverse variety of job profiles were represented, but no software engineers were included. See Table (Table 2) for a full breakdown. Note that we intentionally recruited novice users, as we envision them as intended users of InstructPipe. The criteria for selecting the novice users was based on self-evaluation ratings of the below prompts:
  • Please provide a self-evaluation of your programming experience
  • Please provide a self-evaluation of your machine learning (ML) skillset
Table 2. Participant demographics for the user study, showing various demographic characteristics and skills relevant to InstructPipe.
6.4. Metrics
In addition to the qualitative data from the interview, we measured the following quantitative data.
6.4.1. Task Completion Time
Back-end logs were used to collect timestamps for starting and ending events. Then, completion times for each condition were calculated per task for each participant.
6.4.2. The Number of User Interactions
We used the number of user interactions (introduced in section 5.3) to measure the user’s objective workload. Unlike the results in section 5.4, we report an absolute value here because all the pipelines are controlled in the system evaluation.
6.4.3. Perceived Workload
The raw task load index (Raw-TLX) questionnaire was used to measure participant’s perceived workload (Hart and Staveland, 1988). This questionnaire was a subset of the NASA-TLX (part I). Participants filled out the questionnaire after each condition (InstructPipe or Visual Blocks).
6.5. Results
Figure 11. Raw-TLX results. The statistic significance is annotated with $*{}^{*}$, ${}^{**}$, or ${}^{***}$ (representing $p$¡.05, $p$¡.01, and $p$¡.001, respectively).
Table 3. Task completion time and the number of human interactions in the user study (N=16).
System Time (secs) # Interactions
Median IQR p Median IQR p
InstructPipe 203.5 156.25 p <.001 5.0 4.25 p <.001
Visual Blocks 304.5 124.25 16.0 6.0
6.5.1. InstructPipe Reduces Users’ Workload
[…]
Figure 11 further visualizes the results of users’ perceived workload in six sub-scales. The Wilcoxon signed ranks test revealed significant differences on five sub-scales, all but “Mental Demand” (see section 7.2 for more explanations and discussion). Furthermore, the test indicates that all participants unanimously vote InstructPipe provides lower or equal workload on the subscales of “Physical Demand”, “Temporal Demand”, “Performance” and “Effort” ($W=0$). These quantitative results, with both objective and subjective metrics, demonstrate that InstructPipe can effectively reduce users’ workload during the pipeline creation process.
Users’ qualitative feedback is also aligned with our quantitative results. Participants complimented that InstructPipe is “helpful”[P16] and “obviously easier (to use) than [Visual Blocks]”[P1]. P11 and P6 further elaborated how InstructPipe enhances the user experience when the user builds a visual programming pipeline:
“I feel like I can talk in natural language, and it (InstructPipe) can write the code for me.” [P11]
6.5.2. On-boarding Support of Visual Programming
P1, P5, and P9 explicitly mentioned that there is a “learning curve” in the visual programming system for beginners, which validates our statements in section 1 that motivates this project.
“There is a learning curve to it (using the visual programming system) for sure, because you have to, like, read each node carefully” [P1]
P1’s comment matches with our observation of participants’ behaviors during the study. In the Visual Blocks condition, we observed that people were more easily stuck in their creative purposes, which required our support [^4] In the whole study, we encouraged the users to talk to us while keeping their focus on their task. While we offer explicit hints only when we consider a participant is stuck in an issue, it is still difficult to strictly define what conversation is “ providing hints” and count them quantitatively. Therefore, we cannot fairly report quantitative data for “the number of hints” in both conditions. Here, we faithfully report this finding based on our observation and experience. . Typical supports include 1) guiding participants if we notice they go too far away from the correct pipeline and 2) reminding them of an important node for the pipeline, although we introduced all the necessary nodes in our training session.
To this end, participants commented that InstructPipe is a good onboarding tool in visual programming systems, especially for non-experts, to get familiarized with the system by having a ready solution.
“[InstructPipe] lets you know these nodes exist [when the pipeline appears after the instruction]. It’s like a super speedy tutorial.” [P7]
“If you don’t have experience in visual programming, you will appreciate [InstructPipe] much more … With [InstructPipe], the structure is there, and I feel less worried about making mistakes. It’s, like, giving you examples. It’s easier than starting from scratch.” [P5]
Anecdotally, three participants asked for InstructPipe during the Visual Blocks condition.
6.5.3. Integration into The Existing Workflow
InstructPipe is a feature available in Visual Blocks. In the interviews, participants particularly expressed their strong appreciation of this design as an AI assistant that enhances, instead of completely replacing, the existing user workflow (i.e., the creative process purely using the node-graph editor):
“[The pipeline generated by [InstructPipe] could be pretty close to what I want … Or maybe sometimes not, but that’s okay. I got most of the blocks there, and then it’s up to me to figure out how to connect them.” [P6]
While most participants shared similar opinions with P6, P15 raised one concern about such integration. The concern lies in the mental experience of the state that transits from the prompting task to the node-graph editing process:
“Rather than fixing it (the generated pipeline) on my own [on the node-graph editor], I would have gone back and changed my prompt … I’m pretty sure I could have gotten it closer … because I just spent so much time figuring out what the prompt should be. That’s kind of like already where my brain was and I knew that something was wrong there (the prompt), but I would have to switch over to the other mode (visual programming) of figuring out what was wrong on the pipeline.” [P15]
6.5.4. Use Scenarios: Accessible ML Prototyping and Education
In the open-ended session, we observed participants could efficiently utilize InstructPipe to prototype a pipeline for various daily life or business purposes. For example, P14 tried InstructPipe with “summarize real estate price increase in San Diego California over 2023”. Compared to using LLM chatbots, InstructPipe helps the user quickly build a more explainable pipeline in which the user can track (or modify) the information resources. P4 prototyped an interactive VQA app by “Describe the product in the camera”. P13 further shared his thoughts on how this rapid and accessible prototyping experience can support future business:
“It (InstructPipe)’s going to facilitate the prototype building for PMs (Product Managers) … I was a PM … Back then … My biggest fear is to code … I have lots of ideas, but my challenge is how to translate an idea into the technical world and see a prototype. I think that this app expedites me in that process a lot.” [P13]
Another emerging theme was regarding educating kids on programming as explained in the following:
“With [InstructPipe], I don’t need to teach them (kids) to code for them to build something … Some kids like to code, some kids like to create stuff but don’t want to be bored with learning the syntax of coding … Using [InstructPipe], I can see kids can build, like, customized chat-bots or interactive Wikipedia.” [P13]
6.5.5. Limitations and Future Directions
Across the study sessions, we consistently observed a specific user behavior pattern: participants typically paused their pace when a generated pipeline appeared in the workspace. At these times, some participants used soliloquy, as in saying “Let me see”, while others kept a focused stare on the workspace. These human behaviors strongly indicate that the participants were engaging in deep, contemplative thought.
This observation suggests that participants typically need to perceive a pipeline that appears in the workspace. Such a sense-making process brings new challenges to their creative process:
“[Using InstructPipe] is a little mentally demanding … I have to debug … If it doesn’t help (generating an almost 100% correct pipeline), I have to go through all the nodes … I don’t like debugging.” [P13]
Additionally, we observed that several participants spent more time crafting their prompts than the others. P15 is the participant who spent the most time writing the prompt. The following comments explain the reason, and how the prompting process causes extra mental workload to him/her/them.
“I’m a relatively visual thinker … Getting the prompt right requires me to think in a way that is a lot more like precise and like getting it figured out without working it out live … [When writing prompts, ] you’re just putting them (every detail in a whole pipeline) all out [in one short prompt]” [P15]
In addition to the lack of the original visual thinking experience in visual programming, P13 also warned that such simplification of the creative process into prompting experience may sacrifice users’ hands-on experiences:
“I’m very hands-on with techs. I would like to understand what’s going on [rather than prompting LLMs to generate everything for me]. I want to like think for myself and then compile all the information myself.” [P13]
7. Discussion and Future Directions
7.1. Human-AI Collaboration in Prototyping Open-ended ML Pipelines
Our technical evaluation (section 5.4) shows that InstructPipe reduces the number of user interactions to 18.9 % (±$\pm 20.3\%$). There are two implications from the results.
  • InstructPipe automates most pipeline components with a single prompt.
  • InstructPipe is not able to automate all the pipeline creation processes.
Such takeaways differ from existing findings that show LLMs can achieve full automation of ML inference (Surís et al., 2023; Gupta and Kembhavi, 2023). The main reason is that existing work built their ad hoc solutions for target use scenarios, respectively. In contrast, InstructPipe covers a wide range of mainstream ML models (section 3.2) and aims for an open-ended use case. Our results show that LLMs (i.e., GPT-3.5-turbo) still fail to write robust code with prompt engineering techniques. As we mentioned in section 4.4, the main attribute of such failure is that LLMs tend to invent nodes (i.e., the node type in the pseudocode) that do not exist in our node library, which causes execution error. Note that we explicitly instructed the LLM that “introducing new nodes is not permitted” in the system prompt for the Code Writer module (Figure 7). We also observed similar issues when we combined all the official prompts of four scenarios in VisProg (Gupta and Kembhavi, 2023).
While the execution error in existing work indicates the failure of the full solution, InstructPipe renders the executable components of the generation pipeline in a node-graph editor. This facilitates a collaboration protocol between humans and AI in which humans do not need to start a project from scratch. Results in section 6.5.3 revealed that a wide range of participants appreciated this design of human-AI collaboration. We believe the main reason behind this observation is that most non-expert users prefer working on a task from a template (although sometimes it is not perfect) rather than starting the whole project from scratch. The visual programming system provides non-experts with a novice-friendly interface to work with AI in a visual platform. In this manner, “users can do what matters most – building software by letting AI do the redundant work” (Zhang et al., 2023a).
7.2. Human-AI Collaboration in Prototyping Open-ended ML Pipelines
While most participants are positive about our integration of InstructPipe into the existing user workflow of Visual Blocks, results from the user evaluation also reveal several challenges in such human-AI collaboration. Results in section 6.5.1 show that InstructPipe failed to significantly reduce novice users’ mental demand. We summarized its major causes into three aspects.
  • Instruction. P15’s comment in section 6.5.5 summarizes the first aspect that causes mental burden. Although the “instruction-to-pipeline” process is fast and seems effortless, the process of framing a prompt is one factor that may overwhelm users, especially those who are more accustomed to visual thinking. In essence, an instructor of an LLM system (the user of InstructPipe) should be clear about the problem they want to solve and preferably what pipeline they want, which causes a mental burden to the user. Zamfirescu-Pereira et al.’s study shows that non-experts may not prompt LLMs well (Zamfirescu-Pereira et al., 2023), which further explains why non-expert participants might find the instruction process mentally demanding.
  • Perception. The integration of LLM supports into the visual programming interface enables a “multimodal programming” experience (Dietz et al., 2023), in which, users can program through both verbal and visual approaches. Despite the enhanced flexibility of the system, it also causes additional perceptual burden because users need to switch their “brain mode” between “visual thinking” and “text-based thinking”. This finding is aligned with Dual Coding Theory (DCT) (Paivio, 1991), in which Paivio hypothesized that human brains process information using two different channels: verbal and visual. Interestingly, our results seemingly contradict with multiple psychological findings based on DCT that show a combination of verbal and visual information actually helps humans’ memory process (Paivio, 1969; Paivio et al., 2006). For example, people feel it easier to remember a new word if they learn the word using a vocabulary card with a figure with texts. Based on these existing theories on humans’ mind processes, we make the following explanation on the finding in our project: the cause of mental workload does not imply that people dislike a multimodal workspace, but that users require a transparent interface that connects humans’ mental model to the AI reasoning process, both verbally and visually. When users frame the instruction in the InstructPipe dialogue, there is a lack of feedback system from the AI assistant that visualizes how AI interprets the instructions in real time. As a result, humans need to either first estimate the AI prediction or make no expectations when they perform instruction. Even for experts, it is hard to accurately predict AI generation, and thus, the generated pipeline is usually unexpected. When users frame the instructions, the current interface design is purely based on texts, and the visual channel in the brain is off. When the generation is complete, the user suddenly needs to switch on a sleeping (visual) channel, which causes a mental workload. One possible approach to addressing this issue is to expose users to the multimodal information when they perform instructions. For example, future work can investigate methods to visualize a generated pipeline on-the-fly when users are framing and typing their instructions. In this manner, users’ mental models are continuously synchronized with the AI assistant before the final generation, which should eliminate users’ cognitive burdens when the final generation result appears.
  • Debugging. When a rendered pipeline does not match with users’ expectations, users then are required to debug the generated results (see P13’s comment section 6.5.5). In essence, debugging is one professional programming skill, which explains why it mentally overwhelms beginner-level users. InstructPipe made the first step to facilitate users to 1) perceive the results and 2) take actions on the results, i.e., two important elements in a debugging process, on the node-graph editor. Results in our study reveal that this first step is insufficient to support users so that users still feel like they are debugging the codes as professional programmers. Future work should investigate how to further support users’ debugging process with visual programming interfaces. For example, when execution fails, future work can consider visualizing messages that provide cues on the possible issues and provide actionable guidance on the interface for users to fix the issues easily.
On the other hand, it is important to note that we distill the attributes mainly from a subset of participants who explicitly mentioned that they realized an increase in mental workload when using InstructPipe. Those participants typically spent more time writing prompts and working on the node-graph editor than the average performance (e.g., P15 spent the most time ideating and writing the instructions). Therefore, we argue that the aforementioned issues do not represent the average user experience, and the issues are distilled from explicit comments on their negative experience. In fact, we observed that most participants were relatively decisive and could quickly write their instructions and modify a generated result in the node-graph editor. This observation aligns with our qualitative results in Figure 11: the median score of “Mental Demand” in the InstructPipe condition is lower than the baseline condition.
7.3. Instructing LLMs Poses Challenges for Both Novices and, Potentially, Experts
As we explained above, one reason that caused mental workload is that non-experts found it challenging to instruct LLM. More interestingly, we found that even we, the inventors of InstructPipe, failed to write optimal instructions. As mentioned in section 5.2, two authors annotated captions for the pipelines offline, and we observed multiple imperfections, especially for the complex ones. For instance, the two captions of 8(c) are “Describe the image and turn it into a cat image” and “Edit an image by updating the image caption”. Neither caption explicitly describes the detailed pipeline flow clearly, and therefore, all the six evaluation trials (section 5) were incomplete (see 11(a) for one example). The average ratio of user interactions is $45.8\%$, more than twice the average value for our multimodal pipelines ($20.8\%$). To further understand the cause of the failure, another author improved the instruction into “Caption a tiger image using VQA, modify the character in the caption into a cat using LLM, and finally generate a cat image based on the updated caption”.
((a))
((b))
Figure 12. A comparison of InstructPipe generated by two instructions: (a) “Edit an image by updating the image caption”; (b) “Caption a tiger image using VQA, modify the character in the caption into a cat using LLM, and finally generate a cat image based on the updated caption”. See 8(c) for the complete pipeline.
The resulting pipeline is significantly improved but still not perfect (11(b)). The user only needs to turn “Imagen” into another mode so that it also accepts the input “image” node. Revisiting the improved instruction, we instructed InstructPipe with “generate a cat image based on THE updated caption”, which actually missed the input image.
The important takeaway is while natural languages are proven to be one promising communication media that connects humans and AI systems (Chen et al., 2023; Wang et al., 2023b), instructions may not be the best format to facilitate such connection. We believe the reason is that instructions are still not intuitive to humans: AI typically requires flawless and unambiguous instructions while humans tend to express their intention using ambiguous natural languages in conversations. We encourage future work to investigate alternative interaction mediums beyond instructions to further enhance user experience in human-AI collaboration.
7.4. Online InstructPipe
In this project, we made the following assumption: visual programming and its LLM support are both offline. Note that InstructPipe does have connection to the Internet, and the “offline” here implies that the system does not dynamically update its node library with online resources: every node is pre-defined by the system developers. This statement shows one major limitation of InstructPipe we observed in the open-ended session of our user evaluation: the 27 nodes covered by InstructPipe are insufficient to facilitate all the creative instructions from participants. For example, P9 prompted InstructPipe using “giving me [an] image person walking [a] dog with webcam of myself”. After some clarification, we found that P9 wanted to generate a video in which P9 is walking a dog by taking the webcam as the input of P10’s visual appearance. Such a pipeline is infeasible in our system because it requires a text-driven video generation model. While an extension of InstructPipe with such a text-to-video model as a primitive node can solve this specific issue, we argue that, in practice, users may request many more nodes for their customized uses. Therefore, it is rather important for researchers to investigate a generalizable solution for issues of this category.
In this project, InstructPipe enables users to prototype an ML pipeline with human instruction, and we focus our exploration on a fixed node library. An extension of InstructPipe to a system that covers a wide range of possible online functions is still under-explored. With a next generation of InstructPipe that can find ML models, define nodes and implement the node dynamically with human instruction, we believe the aforementioned issues (e.g., the one raised by P10) can be effectively addressed. In the past several years, the community has already established ML libraries and API services for various models (e.g., by Hugging Face (HuggingFace, 2022)). Such ecosystems provide crucial resources for AI-oriented visual programming systems to dynamically define their primitive nodes, and thus provide more powerful support for users. We encourage future work to investigate how to build a model selector (similar to the Node Selector in Figure 4) that can intelligently select the correct online API to be called in a node. We believe such “online InstructPipe” would provide an unprecedented user experience in which researchers can brainstorm with the system, and the system will automatically return a pipeline with state-of-the-art ML models to accelerate researchers’ creative workflow.
8. Limitations
As discussed in the main content, we notice multiple limitations of InstructPipe. In our current implementation, InstructPipe focuses on generating the graph structure in the pipeline (section 4.1), and therefore the system is not able to generate the property value of the nodes. In some nodes like “tf-model-runner” that exist in Visual Blocks (but not covered by InstructPipe), such property value plays a critical role that directly defines the functionality of the node. Additionally, when InstructPipe detects an undefined generated node, it directly disposes of the line without leveraging such information for other purposes (section 4.4). In the evaluation, we were not able to conduct a large-scale user study with thousands of user study sessions to understand the performance of InstructPipe with both variations from the pipeline factors and the human factors (section 5). Our quantitative results show that InstructPipe fails to significantly reduce participants’ mental workload (fig. 4). Qualitative results further validate this finding and reveal several issues that cause the mental burden when users use InstructPipe (section 7.2).
We also observe two major limitations of the project in addition to those discussed in the main content of the paper. First, our user study participants only engaged InstructPipe in one hour, and therefore, it is unclear whether users will still frequently use and appreciate InstructPipe supports or similar AI assistants in general in a long-term manner. Additionally, all participants are non-experts. Hence, we cannot verify the effectiveness of InstructPipe for users with other levels of expertise. Second, InstructPipe currently cannot detect harmful data or misuse of AI. We believe such a feature is crucial, especially when future work builds “online InstructPipe”(section 7.4), which greatly enhances the generalizability of ML pipeline prototyping capability. Future work must study effective methods to eliminate potential harmful uses when AI assistants become increasingly open-ended.
9. Conclusion
In this paper, we introduce InstructPipe, an AI assistant that facilitates users to build ML visual programming pipelines with text instructions. We design and implement InstructPipe by decomposing the task into three modules: a node selection module, a code writer, and a code compiler. Results in our technical and system evaluations demonstrate that InstructPipe provides users’ satisfactory “on-boarding” experience of visual programming systems and allows them to rapidly prototype an idea. We further discuss the issues we observed concerning LLMs in visual programming, related to both human factors and technical implementations. We hope that InstructPipe will engage a diverse community to easily develop creative machine learning pipelines.
Appendix
Appendix A A Full List of 27 Nodes in InstructPipe
The following content shows 27 nodes InstructPipe covers in the generation process and their corresponding short description used in the Node Selector (section 4.2):
A.1. Input Nodes
  1. (1)
    live_camera: Capture video stream through your device camera.
  2. (2)
    input_image: Select images to use as input to your pipeline. You can also upload your own images.
  3. (3)
    input_text: Add text to use as input to your pipeline.
A.2. Output Nodes
  1. (1)
    image_viewer: View images.
  2. (2)
    threed_photo: Create a 3D photo effect from depthmap tensors.
  3. (3)
    markdown_viewer: Render Markdown strings into stylized HTML.
  4. (4)
    html_viewer: Show HTML content with styles
A.3. Processor Nodes
  1. (1)
    google_search: Use Google to search the web that returns a list of URLs based on a given keyword; usually selected with string_picker.
  2. (2)
    body_segmentation: Segment out people in images; usually selected with mask_visualizer.
  3. (3)
    tensor_to_depthmap: Display tensor data as a depth map.
  4. (4)
    portrait_depth: Generate a 3D depth map for an image; usually selected with tensor_to_depthmap, threed_photo.
  5. (5)
    face_landmark: Detect faces in images. Each face contains 468 keypoints; usually selected with landmark_visualizer, virtual_sticker.
  6. (6)
    pose_landmark: Generate body positional mappings for people detected in images; usually selected with landmark_visualizer.
  7. (7)
    image_processor: Process an image (crop, resize, shear, rotate, change brightness or contrast, add blur or noise).
  8. (8)
    text_processor: Reformat and combine multiple text inputs.
  9. (9)
    mask_visualizer: Visualize masks.
  10. (10)
    string_picker: Select one string from a list of strings; usually used with google_search.
  11. (11)
    image_mixer: Combine images and text into one output image. Requires two image inputs.
  12. (12)
    virtual_sticker: Use face landmarks data to overlay virtual stickers on images.
  13. (13)
    palm_textgen: Generate Text using a large language model.
  14. (14)
    keywords_to_image: Search for images by keywords.
  15. (15)
    url_to_html: Crawl the website by a given URL.
  16. (16)
    image_to_text: Extract text from an image using OCR service.
  17. (17)
    pali: Answer questions about an image using a vision-language model.
  18. (18)
    palm_model: Generate text using a large language model based on prompt and context.
  19. (19)
    imagen: Generate an image based on a text prompt.
  20. (20)
    input_sheet: Read string data from Google Sheets.
((a)) Body segmentation
((b)) PaLI
Figure 13. Examples of node configuration used in Code Writer. The configuration is structured in a JSON format.
Appendix B System Prompts Used in LLM Modules
Here we provide more details about the prompts we utilized in InstructPipe. The original txt files are also attached in the supplementary zip file.
B.1. Node Selector
Please see our supplementary file (node_select.txt) for the full prompt we use in this stage.
B.2. Code Writer
The prompt in Code Writer is dynamic, which is dependent on the nodes chosen in Node Selector. Therefore, we cannot provide all the possible prompts in the supplementary materials. Here, we will focus on providing examples of two detailed node configurations utilized in InstructPipe. Figure 7 shows the structure of the prompt utilized in this LLM stage. Figure 13 provides two examples of node configurations (i.e., “Body segmentation” and “PaLI”) that InstructPipe may chose into the highlighted line(s). Each configuration includes keys of “nodeSpecId” (i.e., node types), “description”, “category” and “examples”. For those nodes that support input and output edges, “inputSpecs” and “outputSpecs” specify the sockets and their corresponding valid data types. For example, the output socket name of “Body segmentation” is “segmentationResult”, and its data type is “masks”. Some nodes (e.g., “Body segmentation”) include recommended node(s) (e.g., “Mask visualizer” for “Body segmentation”), and our configuration also contains such information in the dictionary.
Appendix C User Study Pipelines
Figure 14. Text-based pipeline. The “String picker” node provides users a drop-down menus to select one URL from a list of URLs returned by “Google Search”. “PaLM Text Generator” is an LLM used to summarize the full HTML page.
Figure 15. Real-time multimodal pipeline. The “Keyword to image” node is used to search sunglasses image, and the “Virtual sticker” node anchors the sunglasses onto users’ face.
Figure 14 and Figure 15 visualize two pipelines we required the participants to complete in our user study. Figure 15 is a multimodal pipeline that allows participants to interact with AR effects in real time. Our technical evaluation shows that InstructPipe can generate this pipeline accurately: the averaged ratio of human interactions = 5.2%. Figure 15 is a text-based pipeline that provides participants with a summary of the news searched from Google. The technical evaluation reveals that InstructPipe cannot generate this pipeline accurately: the averaged ratio of human interactions = 27.8%.
Note that even though InstructPipe may be able to complete the pipeline structure in Figure 15 from users’ instruction, we observed that participants still need to fine-tune their keywords to get an ideal pair of sunglasses. Additionally, the default anchor value is “Face top”, so participants need to use the drop-down menu on the “Virtual sticker” node to change the value to “Eyes”. This further motivates us to use the metric of “Time” in addition to the number of user interactions in our study. Our demo video also covers the workflows of these two pipelines.
Appendix D Semi-structured Interview Script
[ Introduction ] ( Start timing! 60 min max. )
Hello, my name is X.
First, I would like to thank you for your participation and completing the consent form. Today, you will be a participant in a user study regarding machine learning and visual programming. Your data will be kept anonymous. Additionally, as a researcher I have no position on this topic and ask that you be as open, honest, and detailed in your answers as possible. Do you have any questions before we begin?
Basically, visual programming borrows the metaphor of block building and allows novice users to develop digital functionalities without writing codes.
[Show Visual Blocks]
Here, each block is called a node, and each node takes in specific inputs, then returns the desired outputs. What you can do is to connect a series of nodes together as a pipeline to achieve a high-level goal.
We are going to walk you through our Visual Blocks system and ask you to actually use Visual Blocks in two conditions to create a few applications.
[ Tutorial ] ( Start timing! 10 min max. )
Before we get started, let us do a tutorial of our system.
[ Study and TLX ] ( Start timing! about 30 min )
[Leverage the counter-balanced sheet and give user a task]
[Think aloud. Have a short discussion with the user. What’s the user’s plan to achieve this given functionality?]
[ Interview ] ( Start timing! about 15 min )
1. What’s your impression of Visual Blocks / InstructPipe [counterbalanced]? Do you need many edits / operations to make it work?
2. Are there any pipelines you come up with in work scenarios / casual scenarios?
3. What works with InstructPipe? In what specific scenarios will InstructPipe be very helpful?
4. What does not work with InstructPipe? Would you give me an example?
5. Do you have any suggestions to improve the design of both systems?
6. Which kinds of technologies would be interesting to add?
7. What applications do you want to start with InstructPipe? And what applications do you want start without it?
That’s all for our user study. Thank you for your participation and we will compensate for your time.